OpenStudy (anonymous):

Find the coordinates of the reflection: B (5, 4) is reflected over the line x=1 B’=? Question 10 options: a) (6, 4) b) (6, 5) c) (-3, 4) d) (4, 4)

OpenStudy (anonymous):

use the same formula but in this case since x=1 is given so x= 2times 1-x ,y be the answer

OpenStudy (anonymous):

sorry for delay .some technical prob occured !

OpenStudy (anonymous):

oh its ok i just don't understand the formula i've never done this before

OpenStudy (anonymous):

this one also c !

OpenStudy (anonymous):

ok thanks :DD

OpenStudy (anonymous):

ok i will give u table so that it'll be easier for u to do other too ok !

OpenStudy (anonymous):

and i have a small question but what is an isometry?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

well isometries are reflection , rotation , translation all thse stuffs

OpenStudy (anonymous):

yes i know that but isnt it a shape that is the same that gets bigger or smaller?

OpenStudy (anonymous):

but streach and shrink are notisometries

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

yah !

OpenStudy (anonymous):

ok so if it stays the same as the original shape its not an isometry?

OpenStudy (anonymous):

for the reflection abt x axis \[x^l=x , y^l= -y \]

OpenStudy (anonymous):

for the reflection abt y axis \[x^l= -x , y^l=y \]

OpenStudy (anonymous):

for the reflection abt y=x \[x^l=x, y^l=y\]

OpenStudy (anonymous):

for the reflection abt y=-x \[x^l=-x, y^l=-y\]

OpenStudy (anonymous):

for the reflection abt x=h where h=1,2,3......\[x^l=2h-x , y^l=y\]

OpenStudy (anonymous):

for the reflection abt y=k where k=1,2,3..... \[x^l=x , y^l= 2k-y \]

OpenStudy (anonymous):

now remember few things then you can easily solve others

OpenStudy (anonymous):

reflection siply means mirror images ! like